Please Help me I never was good at word problem. Please show all work.
Thanks in advance

The length of a rectangular garden is 7 feet longer than its width. If the garden's perimeter is 202 feet, what is the area of the garden in square feet?

Answer by richard1234(7193)

You just have to be able to translate it into an algebra problem.

"The length of a rectangular garden is 7 feet longer than its width"
This means that L = W+7, where L and W are the length and width, in ft.

"If the garden's perimeter is 202 feet"
So 2L + 2W = 202.

"what is the area of the garden in square feet?"
This asks you to find L*W.

So L = W+7, so substituting that into the perimeter equation, we get 2(W+7) + 2W = 202. Can you solve for W and L?